OpenStudy (anonymous):

square root 45+ square root 20 - square root 5

OpenStudy (thread252):

8.94427191

OpenStudy (johnweldon1993):

\[\large \sqrt{45} + \sqrt{20} - \sqrt{5}\] \[\large \sqrt{9 \times 5} + \sqrt{4 \times 5} - \sqrt{5}\] \[\large 3\sqrt{5} + 2\sqrt{5} - \sqrt{5}\] \[\large (3 + 2 - 1)\sqrt{5} = ?\]

OpenStudy (anonymous):

is that the answer

OpenStudy (johnweldon1993):

You have 1 more step to do... \[\large \color \red{(3 + 2 - 1)}\sqrt{5} = ?\]